# Leveraging Discrete Choice Experiments for User-Centric Requirements Prioritization in mHealth Applications

### v1

@@ Abstract
Context: Mobile health (mHealth) applications are widely used for chronic disease management, but usability and accessibility challenges persist due to the diverse needs of users. Adaptive User Interfaces (AUIs) offer a personalized solution to enhance user experience, yet barriers to adoption remain. Understanding user preferences and trade-offs is essential to ensure widespread acceptance of adaptation designs.
@@
Method: A Discrete Choice Experiment (DCE) was conducted with 186 participants who have chronic diseases and use mHealth applications. Participants were asked to select preferred adaptation designs from choices featuring six attributes with varying levels. A mixed logit model was used to analyze preference heterogeneity and determine the factors most likely influencing adoption. Additionally, subgroup analyses were performed to explore differences by age, gender, health conditions, and coping mechanisms.
Results: Maintaining usability while ensuring controllability over adaptations, infrequent adaptations, and small-scale changes are key factors that facilitate the adoption of adaptive mHealth app designs. In contrast, frequently used functions and caregiver involvement can diminish the perceived value of such adaptations.
Conclusion: This study employs a data-driven approach to quantify user preferences, identify key trade-offs, and reveal variations across demographic and behavioral subgroups through preference heterogeneity modeling. Furthermore, our results offer valuable guidance for developing future adaptive mHealth applications and lay the groundwork for continued exploration into requirements prioritization within the field of software engineering.
@@ 1 Introduction
Chronic diseases such as diabetes, heart disease, stroke, and cancer remain leading causes of death worldwide. In 2022, they accounted for nearly 40% of all U.S. deaths and significantly increased COVID-19 risks [hacker2024burden, ahmad2023provisional, People_2024]. Effective chronic disease management relies on sustained adherence to care plans, essential for better health outcomes, quality of life, and cost-efficiency [hamine2015impact]. Mobile health (mHealth) applications provide crucial features like symptom tracking, medication reminders, remote monitoring, and telehealth, which aids in disease management through smartphones and wearables [bricca2022quality]. mHealth applications should enhance accessibility and flexibility to effectively engage users with chronic diseases [choe2017semi]. Most mHealth applications for chronic disease management use a one-size-fits-all design approach that doesn’t meet users’ diverse needs [grua2020reference, reiners2019sociodemographic]. Patients with chronic diseases differ in backgrounds, expertise, and demographic, psychological, and cognitive traits [Vasilyeva2005]. As chronic diseases progress, patients’ needs change, indicating diverse requirements differ not only between individuals but also throughout their lifetime [di2019chronic]. These long-term conditions [harvey2012future] often come with comorbid medical and psychopathological conditions [di2019chronic], increasing user needs and functional requirements [gregor2002designing].
@@
Smartphones and tablets provide flexible access to health information [Vogt2010], while wireless sensors and location systems improve contextual data collection [Vogt2010]. However, these technologies create variability in user environments and needs, making fixed User Interfaces (UIs) inadequate for many mHealth applications, requiring more adaptable solutions to serve diverse users [wang2024adaptive]. Several studies have been conducted on the creation of Adaptive User Interfaces (AUIs) for mHealth applications for various chronic diseases [burke2009optimising, pagiatakis2020intelligent, mohan2008medinet, awada2018adaptive, jabeen2019improving]. Many studies on AUIs have involved minimal user engagement during the development and evaluation phases, making it challenging to assess user preferences for these adaptations [wang2024adaptive]. However, research has consistently shown a trade-off between adaptive mechanisms and usability [paymans2004usability, gajos2008predictability, wesson2010can, lavie2010benefits, peissner2013user, wang2024adaptiveicse, wang2024designing]. These trade-offs complicate UI adaptation decisions, as users differ in prioritizing goals, competence, app usage, and scenarios [wang2024designing]. Limited research has focused on understanding user priorities in adaptive UI design, with existing studies often constrained to specific tasks, applications, and small participant samples [lavie2010benefits, Gajos2017TheIO]. While requirement prioritization is a critical aspect of software development [achimugu2014systematic, riegel2015systematic], traditional approaches tend to focus on stakeholder demands, often overlooking the nuanced user preferences and trade-offs essential for creating truly user-centric designs. Bridging this gap requires systematic methods that capture how users weigh competing design factors in real-world contexts. Discrete Choice Experiments (DCEs) offer a promising yet underexplored approach in Software Engineering (SE) by providing a data-driven, structured method for quantifying user preferences and modeling trade-offs between competing requirements. DCE is a survey-based technique that presents participants with hypothetical choice scenarios to select their preferred option, enabling researchers to infer the relative importance of different attributes and understand how users prioritize competing factors in decision-making [johnson2013constructing]. While DCEs have been widely applied in healthcare and economics to tackle complex decision-making challenges [de2012discrete, tunnessen2020patients, jiang2023patient, al2023patient, jonker2020covid, nittas2020self, simblett2023patient], their potential to enhance requirements prioritization in software development remains largely untapped. This study seeks to address this gap by investigating the key factors influencing AUI adoption in mHealth applications for chronic disease management through DCEs. Our research makes several key contributions to the field of adaptive mHealth design: 1) provides a data-driven approach to user preference quantification and assess the relative importance of key adaptation attributes in mHealth applications; 2) uncovers key trade-offs of competing design attributes and preference heterogeneity across different user subgroups and 3) lays the foundation for developing next-generation adaptive mHealth applications while outlining potential research directions in SE.
The rest of the paper is organized as follows. Section 2 reviews the related work, focusing on the trade-offs between different adaptation factors and the application of DCE in healthcare research and software engineering research. Section 3 outlines the research methodology used to conduct the DCE study. The results are presented in Section 4, followed by Section 5, which explores key implications for mHealth developers and researchers. Sections 6 and 7 conclude by outlining the study’s limitations, contributions, and offering suggestions for future research directions.
@@ 2.1 Adaptive User Interfaces for mHealth applications
AUIs have been deployed across a spectrum of mHealth applications, ranging from stroke rehabilitation [burke2009optimising], diabetes [pagiatakis2020intelligent], cardiac disease [mohan2008medinet], dementia [awada2018adaptive] and Parkinson’s disease [jabeen2019improving]. A systematic literature review (SLR) [wang2024adaptive] identified three primary types of Adaptive User Interfaces (AUIs) in mHealth applications: presentation adaptation, content adaptation, and behavior adaptation. Presentation adaptation focuses on enhancing user experience by modifying the visual aspects of the UI, such as adjusting colors, positioning, and font sizes to better align with user preferences and accessibility needs. Content adaptation tailors the information displayed to better match individual user needs. For example, this can include simplified text representation that enhances readability through clear and concise language (see Figure 1(a)), text-to-image conversion that transforms complex textual content into visual formats for better understanding (see Figure 1(b)), and interface element rearrangement (see Figure 1(c)). Behavior adaptation is a more complex form of adaptation that spans multiple aspects of content and presentation while also modifying system interactions. For example, this can involve adjusting navigation permissions, customizing motivation techniques to encourage behavior change according to user type or health status, and modifying interface modalities to accommodate different usage contexts. These mechanisms enable mHealth apps to deliver personalized and accessible experiences aligned with user needs.
@@ 2.2 Trade-offs between different UI adaptation factors
Existing research on AUIs for mHealth often lacks detailed reporting on the design and development process, particularly in the early stages involving user input, making it challenging to align user preferences and requirements with the final AUI solutions [wang2024adaptive]. Additionally, many evaluations assess the overall effectiveness of applications rather than specific adaptive features, often using small participant sample, which limits the generalizability of findings on AUI effectiveness [wang2024adaptive]. As a result, the limited involvement of users in both the development and evaluation of AUIs raises uncertainty about whether users truly prefer adaptive interfaces as a whole or specific adaptive functionalities within them. Further, research has identified distinct trade-offs between adaptation features and design choices [paymans2004usability, gajos2008predictability, wesson2010can, lavie2010benefits, peissner2013user, wang2024adaptiveicse]. Adaptations can improve or alter usability [paymans2004usability, wesson2010can, wang2024adaptiveicse], and may introduce predictability challenges, making it difficult for users to anticipate subsequent actions [gajos2008predictability]. The trade-offs associated with adaptation are influenced by the balance between routine and non-routine tasks, with older individuals experiencing greater adaptation benefits than their younger counterparts [lavie2010benefits]. peissner2013user investigated the balance between adaptability and user control by analyzing situations that require explicit user confirmation prior to adaptation and those that do not. These studies are restricted by a limited number of participants and a specific demographic profile of users, such as older participants, and focus on specific adaptation conditions. The Technology Acceptance Model (TAM) [davis1989technology] and Unified Theory of Acceptance and Use of Technology (UTAUT) [venkatesh2012consumer] offer a lens for understanding these trade-offs: adaptive features can enhance perceived usefulness and ease of use if they align with user goals, but unpredictability or poor transparency may hinder adoption. UTAUT also highlights the role of facilitating conditions (e.g., customization options). However, it is still unclear how users’ implicit preferences align with these trade-offs, the level of control they expect over system-driven decisions, and how priority variations differ across user groups, as only a few studies have explored these aspects. Although advances in AI allow systems to learn and adapt to human behavior dynamically, the lack of transparency in how such systems prioritize features raises critical questions [ehsan2021expanding]. Understanding these priorities is crucial to ensure that AI systems adapt in ways that align with user expectations, improve usability, and build trust.
@@ 2.3 Requirements prioritization
Requirements prioritization is a crucial aspect of software development, particularly in mobile application development, where frequent system updates, limited resources, response to user feedback, and challenges such as decreased user adherence due to overwhelming choices complicate the process [jantunen2011challenge]. Given the constraints of human resources, technical feasibility, cost, and schedule, it is rarely feasible to implement all the identified requirements. Therefore, the primary goal of requirements prioritization is to select an optimal set of requirements that balances stakeholder demands with available resources while effectively addressing trade-offs [pitangueira2015software]. Although several SLRs have examined prioritization techniques [achimugu2014systematic, riegel2015systematic], most of these approaches focus on meeting the needs of project stakeholders rather than incorporating the preferences and priorities of the end users. This is particularly problematic in user-centric software, where conflicting requirements and trade-offs between usability, functionality, and other factors are common [johnson2020designing].
@@ 2.4 Discrete choice experiments
Stated preference methods are survey-based approaches designed to elicit people’s preferences towards specific, often poorly understood behaviors, with DCE being the most widely used type [johnston2017contemporary, johnson2013constructing]. In recent years, DCEs have become accepted as a useful method for quantifying societal preferences and setting priorities in healthcare by allowing an understanding of the relative importance of attributes affecting healthcare decisions and especially the trade-offs that individuals are willing to make between these attributes [de2012discrete, soekhai2019discrete]. Numerous DCE studies have been reported that address patient experiences and health outcomes [de2012discrete]. Some research explores optimal methods for providing health services or medical treatments. For example, certain studies focus on patients’ treatment preferences for depressive and anxiety disorders [tunnessen2020patients]. A study investigates how patients weigh their options among different attributes of targeted pharmacotherapy, such as cost, effectiveness, and side effects [jiang2023patient], and other studies are dedicated to epilepsy treatment [al2023patient].
@@
The development of mHealth technologies that improve access to health knowledge and information is crucial for addressing physical and social inequities [han2010professional], particularly given the high levels of non-adherence to treatment regimens in developing nations [beaglehole2008improving, alwan2009review]. However, the implementation of mHealth applications presents significant challenges, with mixed results regarding their effectiveness in supporting adherence to treatment plans [hamine2015impact, gandapur2016role]. Another key issue is low adoption and high abandonment rates among users [greenhalgh2017beyond], underscoring a persistent gap between app design and consumer preferences [hamine2015impact]. Few studies have explored strategies for improving mHealth technology design, including COVID-19 contact tracing applications [jonker2020covid] and self-monitoring tools that track sun exposure duration and intensity [nittas2020self]. Additionally, research has examined patient preferences for key factors driving the adoption of mHealth solutions for mental health management, such as those addressing depression [simblett2023patient].
@@
Traditional requirements elicitation methods often fail to capture the complexities of user decision-making and trade-offs. While all requirements are typically considered essential, research shows that their importance varies among stakeholders [achimugu2014systematic]. DCEs offer a solution by simulating real-world decision-making, requiring users to weigh multiple trade-offs simultaneously rather than evaluating features in isolation [johnston2017contemporary, johnson2013constructing]. Unlike standard surveys, which rank or rate attributes independently, DCEs present choice scenarios with varying attribute combinations, enabling researchers to quantify user preferences and support more nuanced decision-making [brett2013quantifying]. Integrating DCE into software development presents a novel opportunity to align mHealth features with user expectations and enhance prioritization effectiveness. However, despite its proven benefits in other fields [de2012discrete], DCEs have yet to be widely adopted in SE, highlighting a critical gap in leveraging this method for informed design and decision-making in this field.
@@ 3 Method
We conducted a DCE survey study to explore the perceptions of individuals with chronic diseases and prior experience using mHealth applications on various aspects of adaptation. In the survey, participants were repeatedly presented with two versions of the same mHealth application and asked to select their preferred option (see example question in Figure 2). Both options shared the same set of attributes, but each varied in attribute levels. By analyzing these repeated choices, we can statistically determine which attributes—and their specific levels—participants value most, providing insights into the factors that maximize user utility [ryan2001use, ryan2000using].
Our methodology follows the guidelines reported by the ISPOR Good Research Practices for Conjoint Analysis Task Force [bridges2011conjoint] and is structured around five steps: (1) Key attributes identification, (2) Survey development, (3) Piloting and survey adjustments, (4) Data collection, and (5) Choice data analysis. The survey was administered online and received ethical approval from the Monash University Human Research Ethics Committee (Project ID: 40245).

### v2

# User-Centric Requirements Prioritization in mHealth Applications: Insights from a Discrete Choice Experiment

@@ Abstract
Context: Mobile health (mHealth) applications are widely used for chronic disease management, but usability and accessibility challenges persist due to the diverse needs of users. Adaptive User Interfaces (AUIs) offer a promising approach to personalizing interactions and improving user experience. However, their adoption remains limited, partly due to a lack of understanding of how users perceive and evaluate different adaptation strategies. Addressing this gap is crucial for advancing user-centered design and requirements engineering in software systems for health contexts.
@@
Method: A Discrete Choice Experiment (DCE) was conducted with 186 participants living with chronic conditions who regularly use mHealth applications. Each participant completed a series of choice tasks, selecting their preferred adaptation designs from scenarios composed of six attributes with varying levels. A mixed logit model was applied to examine preference heterogeneity. Subgroup analyses were also conducted to explore variations in preferences across age, gender, health condition, and coping mechanism.
Results: Participants preferred adaptation designs that preserved usability, offered controllability, introduced changes infrequently, and applied small-scale modifications. Conversely, adaptations affecting frequently used functions and those involving caregiver input were generally viewed less favorably. These findings highlight key trade-offs that influence user acceptance of adaptive mHealth interfaces.
Conclusion: This study employs a data-driven approach to quantify user preferences, identify key trade-offs, and reveal variations across demographic and behavioral subgroups through preference heterogeneity modeling. These insights provide actionable guidance for designing more user-centered adaptive interfaces and contribute to advancing requirements prioritization practices in software engineering—particularly in the context of health technologies.
@@ 1 Introduction
Chronic diseases such as diabetes, heart disease, stroke, and cancer remain leading causes of death worldwide. In 2022, they accounted for nearly 40% of all U.S. deaths and significantly increased COVID-19 risks [38, 2, 29]. Effective chronic disease management relies on sustained adherence to care plans, essential for better health outcomes, quality of life, and cost-efficiency [40]. Mobile health (mHealth) applications provide crucial features like symptom tracking, medication reminders, remote monitoring, and telehealth, which aids in disease management through smartphones and wearables [15]. mHealth applications should enhance accessibility and flexibility to effectively engage users with chronic diseases [22]. Most mHealth applications for chronic disease management use a one-size-fits-all design approach that doesn’t meet users’ diverse needs [37, 75]. Patients with chronic diseases differ in backgrounds, expertise, and demographic, psychological, and cognitive traits [93]. As chronic diseases progress, patients’ needs change, indicating diverse requirements differ not only between individuals but also throughout their lifetime [26]. These long-term conditions [42] often come with comorbid medical and psychopathological conditions [26], increasing user needs and functional requirements [36]. However, few studies have systematically addressed how to design mHealth systems that can adapt to these complex, dynamic, and individualized user needs over time.
@@
Smartphones and tablets provide flexible access to health information [95], while wireless sensors and location systems improve contextual data collection [95]. However, these technologies create variability in user environments and needs, making fixed User Interfaces (UIs) inadequate for many mHealth applications, requiring more adaptable solutions to serve diverse users [97]. Several studies have been conducted on the creation of Adaptive User Interfaces (AUIs) for mHealth applications for various chronic diseases [19, 69, 66, 11, 47]. Many studies on AUIs have involved minimal user engagement during the development and evaluation phases, making it challenging to assess user preferences for these adaptations [97]. However, research has consistently shown a trade-off between adaptive mechanisms and usability [72, 31, 100, 60, 73, 98, 96]. These trade-offs complicate UI adaptation decisions, as users differ in prioritizing goals, competence, app usage, and scenarios [96]. Limited research has focused on understanding user priorities in adaptive UI design, with existing studies often constrained to specific tasks, applications, and small participant samples [60, 30]. While requirement prioritization is a critical aspect of software development [1, 76], traditional approaches tend to focus on stakeholder demands, often overlooking the nuanced user preferences and trade-offs essential for creating truly user-centric designs. Bridging this gap requires systematic methods that capture how users weigh competing design factors in real-world contexts. Our study contributes to this gap by using a large-scale Discrete Choice Experiment to empirically model user trade-offs and preferences in adaptive mHealth UI design across diverse user groups.
Discrete Choice Experiments (DCEs) offer a promising yet underexplored approach in Software Engineering (SE) by providing a data-driven, structured method for quantifying user preferences and modeling trade-offs between competing requirements. DCE is a survey-based technique that presents participants with hypothetical choice scenarios to select their preferred option, allowing researchers to infer the relative importance of different attributes and understand how users prioritize competing factors in decision-making [50]. Although DCEs have been widely applied in healthcare and economics to tackle complex decision-making challenges [25, 91, 49, 6, 54, 67, 84], their potential to improve requirements prioritization in software development remains largely unexplored. This study seeks to address this gap by investigating the key factors influencing AUI adoption in mHealth applications for chronic disease management through DCEs. Our research makes several key contributions to the field of adaptive mHealth design. First, it provides a data-driven approach to quantifying user preferences and assessing the relative importance of key adaptation attributes in mHealth applications. Second, it uncovers important trade-offs between competing design factors and reveals preference heterogeneity across user subgroups based on demographic and behavioral characteristics. Finally, it lays the foundation for the development of next-generation adaptive mHealth applications and identifies future research directions within the broader field of software engineering. The rest of the paper is organized as follows. Section 2 reviews the related work, focusing on the trade-offs between different adaptation factors and the application of DCE in healthcare research and software engineering research. Section 3 outlines the research methodology used to conduct the DCE study. The results are presented in Section 4, followed by Section 5, which explores key implications for mHealth developers and researchers. Sections 6 and 7 conclude by outlining the study’s limitations, contributions, and offering suggestions for future research directions.
@@ 2.1 Adaptive User Interfaces for mHealth applications
AUIs have been deployed across a spectrum of mHealth applications, ranging from stroke rehabilitation [19], diabetes [69], cardiac disease [66], dementia [11] and Parkinson’s disease [47]. A systematic literature review (SLR) [97] identified three primary types of Adaptive User Interfaces (AUIs) in mHealth applications: presentation adaptation, content adaptation, and behavior adaptation. Presentation adaptation focuses on improving user experience by modifying the visual aspects of the UI, such as adjusting colors, positioning, and font sizes to better align with user preferences and accessibility needs. Content adaptation tailors the information displayed to better match individual user needs. For example, this can include simplified text representation that enhances readability through clear and concise language (see Figure 1(a)), text-to-image conversion that transforms complex textual content into visual formats for better understanding (see Figure 1(b)), and interface element rearrangement (see Figure 1(c)). Behavior adaptation is a more complex form of adaptation that spans multiple aspects of content and presentation while also modifying system interactions. For example, this can involve adjusting navigation permissions, customizing motivation techniques to encourage behavior change according to user type or health status, and modifying interface modalities to accommodate different usage contexts. These mechanisms enable mHealth apps to deliver personalized and accessible experiences that align with user needs.
@@ 2.2 Trade-offs between different UI adaptation factors
Existing research on AUI for mHealth often lacks detailed reporting on the design and development process, particularly in the early stages that involve user input, making it difficult to align user preferences and requirements with the final AUI solutions [97]. In addition, many evaluations assess the overall effectiveness of applications rather than specific adaptive features, often using small participant sample, which limits the generalizability of findings on AUI effectiveness [97]. As a result, the limited involvement of users in both the development and evaluation of AUIs raises uncertainty about whether users truly prefer adaptive interfaces as a whole or specific adaptive functionalities within them. In addition, research has identified distinct trade-offs between adaptation features and design choices. Adaptations can improve or alter usability [72, 100, 98], and may introduce predictability challenges, making it difficult for users to anticipate subsequent actions [31]. The trade-offs associated with adaptation are influenced by the balance between routine and non-routine tasks, with older individuals experiencing greater adaptation benefits than their younger counterparts [60]. Peissner and Edlin-White [73] investigated the balance between adaptability and user control by analyzing situations that require explicit user confirmation prior to adaptation and those that do not. These studies are restricted by a limited number of participants and a specific demographic profile of users, such as older participants, and focus on specific adaptation conditions. The Technology Acceptance Model (TAM) [24] and the Unified Theory of Acceptance and Use of Technology (UTAUT) [94] offer a lens for understanding these trade-offs: adaptive features can enhance perceived usefulness and ease of use if they align with user goals, but unpredictability or poor transparency may hinder adoption. UTAUT also highlights the role of facilitating conditions (e.g., customization options). However, it is still unclear how users’ implicit preferences align with these trade-offs, the level of control they expect over system-driven decisions, and how priority variations differ between user groups, as only a few studies have explored these aspects. Although advances in AI allow systems to learn and adapt to human behavior dynamically, the lack of transparency in how such systems prioritize features raises critical questions [27]. Understanding these priorities is crucial to ensure that AI systems adapt in ways that align with user expectations, improve usability, and build trust.
@@ 2.3 Requirements prioritization
Requirements prioritization is a crucial aspect of software development, particularly in mobile application development, where frequent system updates, limited resources, response to user feedback, and challenges such as decreased user adherence due to overwhelming choices complicate the process [48]. Given the constraints of human resources, technical feasibility, cost, and schedule, it is rarely feasible to implement all the identified requirements. Therefore, the primary goal of requirements prioritization is to select an optimal set of requirements that balances stakeholder demands with available resources while effectively addressing trade-offs [74]. Several SLRs have examined requirements prioritization techniques [1, 76]; these works provide a comprehensive overview of existing methods and are complementary to our study. However, most of the approaches they review primarily emphasize the perspectives of project stakeholders rather than directly capture end-user preferences. This limitation is particularly evident in user-centric software systems, where conflicts and trade-offs between usability, functionality, and other quality attributes are common [52]. In contrast, our work foregrounds end-user preference elicitation as a core input to requirements prioritization.
@@ 2.4 Discrete choice experiments
Stated preference methods are survey-based approaches designed to elicit people’s preferences towards specific, often poorly understood behaviors, with DCE being the most widely used type [53, 50]. In recent years, DCEs have become accepted as a useful method for quantifying societal preferences and setting priorities in healthcare by allowing an understanding of the relative importance of attributes affecting healthcare decisions and especially the trade-offs that individuals are willing to make between these attributes [25, 86]. Numerous DCE studies have been reported that address patient experiences and health outcomes [25]. Some research explores optimal methods for providing health services or medical treatments. For example, certain studies focus on patients’ treatment preferences for depressive and anxiety disorders [91]. A study investigates how patients weigh their options among different attributes of targeted pharmacotherapy, such as cost, effectiveness, and side effects [49], and other studies are dedicated to the treatment of epilepsy [6].
@@
The development of mHealth technologies that improve access to health knowledge and information is crucial to address physical and social inequities [41], particularly given the high levels of non-adherence to treatment regimens in developing nations [12, 9]. However, the implementation of mHealth applications presents significant challenges, with mixed results in their effectiveness in supporting the adherence to treatment plans [40, 32]. Another key issue is the low adoption and high abandonment rates among users [35], underscoring a persistent gap between the design and consumer preferences [40]. Few studies have explored strategies for improving mHealth technology design, including COVID-19 contact tracing applications [54] and self-monitoring tools that track sun exposure duration and intensity [67]. Additionally, research has examined patient preferences for key factors driving the adoption of mHealth solutions for mental health management, such as those that address depression [84].
@@
Traditional requirements elicitation methods often fail to capture the complexities of user decision-making and trade-offs. Although all requirements are typically considered essential, research shows that their importance varies between stakeholders [1]. DCEs offer a solution by simulating real-world decision-making, requiring users to weigh multiple trade-offs simultaneously rather than evaluating features in isolation [53, 50]. Unlike standard surveys, which rank or rate attributes independently, DCEs present choice scenarios with varying attribute combinations, enabling researchers to quantify user preferences and support more nuanced decision-making [14]. Integrating DCE into software development presents a novel opportunity to align mHealth features with user expectations and enhance prioritization effectiveness. However, despite its proven benefits in other fields [25], DCEs have not yet been widely adopted in SE, highlighting a critical gap in using this method for informed design and decision-making in this field.
@@ 3 Method
We conducted a DCE survey study to explore the perceptions of individuals with chronic diseases and prior experience using mHealth applications on various aspects of adaptation. Given the need for a focused and interpretable approach in a short communication format, the DCE method offers a concise yet powerful way to quantify user preferences through structured choice tasks. In the survey, participants were repeatedly presented with two versions of the same mHealth application and asked to select their preferred option (see the example question in Figure 2). By analyzing these repeated choices, we can statistically determine which attributes—and their specific levels—participants value most, providing information on factors that maximize user utility [79, 80].
Our methodology follows the guidelines reported by the ISPOR Good Research Practices for Conjoint Analysis Task Force [16] and is structured around five steps: (1) Key attributes identification, (2) Survey development, (3) Piloting and survey adjustments, (4) Data collection, and (5) Choice data analysis. The survey was administered online and received ethical approval from the Monash University Human Research Ethics Committee (Project ID: 40245).